OpenStudy (anonymous):

How can polynomial functions be written when given the zeros? can anybody give me a simple answer no examples

OpenStudy (welshfella):

given the zeros a , b and c we can write P(x) = (x - a)(x - b)(x - c)

OpenStudy (zzr0ck3r):

there will be infinite polynomials with those roots.

OpenStudy (welshfella):

you can the expand this to obtain standard form

OpenStudy (zzr0ck3r):

\(p(x)=k(x-a)(x-b)(x-c)\)

OpenStudy (zzr0ck3r):

and of course, if there were more roots there would be more binomial factors...

OpenStudy (welshfella):

ah yes i missed the constant k

OpenStudy (anonymous):

can you put that all together please CX it is confusing me CX lol

OpenStudy (welshfella):

what part don't you understand - why the k had to be added ?

OpenStudy (welshfella):

k can be any constant:- 2 , 6, 0.25 whatever

OpenStudy (welshfella):

and following on from zzrock's post you can have more than 3 roots - for each root there will be another binomial for 4 roots we would have p(x) = k(x-a)(x-b)(x-c)(x-d)